# Towards diffusion models for large-scale sea-ice modelling

### v1

@@ 1 Introduction
Generative diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) has revolutionised data generation in computer vision (Dhariwal & Nichol, 2021; Saharia et al., 2022) and other domains (Bar-Tal et al., 2024; Huang et al., 2023). Initial applications to geophysical systems show promise for prediction (Leinonen et al., 2023; Li et al., 2023; Price et al., 2024; Finn et al., 2024) and downscaling (Mardani et al., 2023; Wan et al., 2023; Brajard et al., 2024). To cut costs, data can be mapped into a latent space using a pre-trained autoencoder, where diffusion models can be learned (Sinha et al., 2021; Vahdat et al., 2021; Rombach et al., 2022). However, the mapping can reduce the model’s ability to generate fine-grained data (Dai et al., 2023).
Here, we examine how well latent diffusion models (LDMs) can generate geophysical data compared to diffusion directly applied in data space. We use Arctic-wide sea-ice simulations that were run with the state-of-the-art sea-ice model neXtSIM (Rampal et al., 2016; Ólason et al., 2022) at a $\frac{1}{4}^{\circ}$ resolution (Boutin et al., 2023). Sea ice, with its discrete elements, anisotropy, and multifractality, poses a challenging problem for generative models and serves as ideal candidate for evaluating LDMs. We show that while the general structure remains the same, LDMs produces smoother solutions than diffusion models in data space, as shown in Fig. 1.
@@
Although generative diffusion is fundamently defined by its Gaussian diffusion process, we can integrate prior physical information into the encoder and decoder of an LDM, as similarly done in Shmakov et al. (2023). Hence, we tailor the approach of LDMs to Earth system modelling by incorporating physical bounds into the autoencoder. To bound the variables, we replace the common Gaussian reconstruction loss (mean-squared error) by the log-likelihood of censored Gaussian distributions, which combines regression and classification, enabling us to explicitly account for the bounds of the output during reconstruction. Such censored distributions improve the representation of the sea-ice extent and thereby enhance the physical consistency of the output, as demonstrated in this study.
@@ 3 Results
Model $\mathrm{FAED}$downarrow $\text{RMSE}_{\text{SIE}}$downarrow
@@
Diffusion 1.73 9.66
LDM 1.65 11.56
LDM (censored) 1.79 7.32
Inspired by the Fréchet inception distance (Heusel et al., 2017), we evaluate the diffusion models in comparison to the testing dataset with the Fréchet distance in the latent space of an independently trained VAE ($\mathrm{FAED}$). We additionally compare the probability that a grid point is covered by sea ice with the root-mean-squared error $\text{RMSE}_{\text{SIE}}$, with results shown in Table 2. Whereas we evaluate on $4383$ generated samples only, the results are robust to the dataset size as demonstrated in Appendix E.2.
@@
Diffusion models are trained with a Gaussian diffusion process, and the data space model struggles to accurately represent the sea-ice extent, leading to a higher RMSE than for the validation dataset. The smoothing in the LDM exacerbates the inaccuracies in representing sea-ice extent, resulting in even higher RMSE values. However, as for the reconstruction, censoring improves the sea-ice extent representation, bringing the RMSE to the realm of the validation dataset.
@@ 4 Conclusions
In this manuscript, we make the first step towards using latent diffusion models (LDMs) for generating multivariate, Arctic-wide sea-ice data. We compare LDMs with diffusion models in data space and find that LDMs lose small-scale information compared to the target fields, though their scores are similar to those of diffusion models in data space. This loss of information is due to the latent mapping process and its pre-training as variational autoencoder. To mitigate this, we could increase the number of channels in latent space (see Appendix E.1). Another approach would be to pre-train the autoencoder with an additional adversarial (Rombach et al., 2022) or spectral loss (Kochkov et al., 2023), although these methods could complicate and destabilise the training and tuning process.
Despite the loss of small-scale information, LDMs offer a significant advantage over diffusion models in data space: the ability to enforce physical bounds during data generation. By replacing the Gaussian data log-likelihood by a censored log-likelihood, we explicitly incorporate the clipping of the decoder output into the training process. The censoring has a neutral impact on the sample quality but leads to a more accurate representation of the sea-ice extent, indicating an increased physical consistency. This increased consistency could benefit other tasks as surrogate modelling or downscaling.
@@ Acknowledgments
We present here work whose goal is to advance the field of Earth system modeling and machine learning. While there are many potential societal consequences of our work, we will briefly discuss the impact on the energy consumption only.
@@ Appendix B Architectures
In the following, we will describe our neural network architectures. The autoencoder is purely based on a convolutional architecture, while the diffusion models are based on a mixture between a convolutional U-Net (Ronneberger et al., 2015) and a vision transformer (ViT, Dosovitskiy et al. (2021)), termed U-ViT (Hoogeboom et al., 2023).
@@ B.3 Diffusion architecture
The diffusion models modify the U-ViT architectures as introduced in Hoogeboom et al. (2023). The encoding and decoding part of the architecture are implemented with convolutional operations, similar to how we implemented the autoencoder. In the bottleneck, the architecture uses a vision transformer (Dosovitskiy et al., 2021).
@@ Appendix C Experiments
Given these datasets, we optimise all models with Adam (Kingma & Ba, 2017) with a batch size of 16. The learning rate is linearly increased from $1\times 10^{-6}$ to $2\times 10^{-4}$ within $5\times 10^{3}$ iterations. Afterwards a cosine scheduler is used to decrease the learning rate again up to $1\times 10^{-6}$ after $10^{5}$ iterations, where the training is stopped. All models are trained without weight decay and other regularisation techniques like dropout, and we select the models that achieve the lowest validation loss. We train the models with mixed precision in bfloat16 and evaluate in single precision. All models are trained on either a NVIDIA RTX5000 GPU (24 GB) or a NVIDIA RTX6000 GPU (48 GB); to train on the RTX5000, we use a batch size of 8 and accumulate the gradient of two iterations. All models are implemented in PyTorch (Paszke et al., 2019) with PyTorch lightning (Falcon et al., 2020) and Hydra (Yadan, 2019). The code will be published for the camera-ready version of the paper.
@@
Both, the reconstructions and the generated data are compared to the testing dataset. For the reconstructions, we perform a one-to-one comparison, while we evaluate the statistics of our generated data. In total, we hence generate $4383$ samples with the diffusion models. Although a smaller number than the $50000$ samples normally used in validating image diffusion models, it results into a fair comparison to the testing dataset. In Appendix, we demonstrate that the dataset size is big enough to see differences in the models.
@@ E.2 Influence of the generated dataset size on the diffusion scores
E.3 Censoring
@@
E.4 Uncurated data samples

### v2

@@ 1 Introduction
Generative diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) has revolutionised data generation in computer vision (Dhariwal & Nichol, 2021; Saharia et al., 2022a) and other domains (Bar-Tal et al., 2024; Huang et al., 2023). Initial applications to geophysical systems show promise for prediction (Leinonen et al., 2023; Li et al., 2023; Price et al., 2024; Finn et al., 2024) and downscaling (Mardani et al., 2023; Wan et al., 2023; Brajard et al., 2024). To cut costs, data can be mapped into a latent space using a pre-trained autoencoder, where diffusion models can be learned (Sinha et al., 2021; Vahdat et al., 2021; Rombach et al., 2022). However, the mapping can reduce the model’s ability to generate fine-grained data (Dai et al., 2023).
Here, we examine how well latent diffusion models (LDMs) can generate geophysical data compared to diffusion directly applied in data space. As a proof-of-concept, we unconditionally generate Arctic-wide sea-ice states as simulated with the state-of-the-art sea-ice model neXtSIM (Rampal et al., 2016; Ólason et al., 2022) at a $\frac{1}{4}^{\circ}$ resolution (Boutin et al., 2023). Sea ice, with its discrete elements, anisotropy, and multifractality, poses a challenging problem for generative models and serves as ideal candidate for evaluating LDMs. We show that while the general structure remains the same, LDMs produces smoother solutions than diffusion models in data space, as presented in Fig. 1.
@@
Although generative diffusion is fundamentally defined by its Gaussian diffusion process, we can integrate prior physical information into the encoder and decoder of an LDM, as similarly done in Shmakov et al. (2023). Hence, we tailor the approach of LDMs to Earth system modelling by incorporating physical bounds into the autoencoder. To bound the variables, we replace the common Gaussian reconstruction loss (mean-squared error) by the log-likelihood of censored Gaussian distributions, which combines regression and classification, enabling us to explicitly account for the bounds of the output during reconstruction. Such censored distributions improve the representation of the sea-ice extent and thereby enhance the physical consistency of the output, as demonstrated in this study.
@@ 3 Results
[…]
@@
Diffusion 1.73 9.66 1.764
LDM 1.65 11.56 0.069
LDM (censored) 1.79 7.32 0.070
Inspired by the Fréchet inception distance (Heusel et al., 2017), we evaluate the diffusion models in comparison to the testing dataset with the Fréchet distance in the latent space of an independently trained VAE ($\mathrm{FAED}$). We additionally compare the probability that a grid point is covered by sea ice with the root-mean-squared error $\text{RMSE}_{\text{SIE}}$, with results shown in Table 2. Whereas we evaluate on $4383$ generated samples only, the results are robust to the dataset size as demonstrated in Appendix E.2. Additionally, we measure the speed (latency) of the diffusion models in seconds for a single generation.
@@
Diffusion models are trained with a Gaussian diffusion process, and the data space model struggles to accurately represent the sea-ice extent, leading to a higher RMSE than for the validation dataset. The smoothing in the LDM exacerbates the inaccuracies in representing sea-ice extent, resulting in even higher RMSE values. However, as for the reconstruction, censoring improves the sea-ice extent representation, bringing the RMSE to the realm of the validation dataset. In addition to an improved representation, the latent diffusion models are 25×25$\times$25 × faster than the diffusion model in data space.
@@ 4 Conclusions
In this manuscript, we make the first step towards using latent diffusion models (LDMs) for generating multivariate, Arctic-wide sea-ice data. We compare LDMs with diffusion models in data space and find that LDMs lose small-scale information compared to the target fields, though their scores are similar to those of diffusion models in data space. This loss of information is due to the latent mapping process and its pre-training as variational autoencoder. To mitigate this, we could increase the number of channels in latent space (see Appendix E.1). Another approach would be to pre-train the autoencoder with an additional adversarial (Rombach et al., 2022) or spectral loss (Kochkov et al., 2023). These methods to try to mitigate the smoothing could complicate (e.g., Rombach et al., 2022) and notoriously destabilise (e.g., Karras et al., 2018) the training and tuning process.
Despite the loss of small-scale information, LDMs offer a significant advantage over diffusion models in data space: the ability to enforce physical bounds during data generation and their speed. By replacing the Gaussian data log-likelihood by a censored log-likelihood, we explicitly incorporate the clipping of the decoder output into the training process. The censoring has a neutral impact on the sample quality but leads to a more accurate representation of the sea-ice extent, indicating an increased physical consistency. This increased consistency could benefit other tasks as surrogate modelling or downscaling.
@@ Acknowledgments
We present here work whose goal is to advance the field of Earth system modelling and machine learning. While there are many potential societal consequences of our work, we will briefly discuss the impact on the energy consumption only.
@@ Appendix B Architectures
In the following, we will describe our neural network architectures. The autoencoder is purely based on a convolutional architecture, while the diffusion models are based on a mixture between a convolutional U-Net (Ronneberger et al., 2015) and a vision transformer (ViT, Dosovitskiy et al. (2021)), termed UVit (Hoogeboom et al., 2023).
@@ B.3 Diffusion architecture
The diffusion models modify the UViT architectures as introduced in Hoogeboom et al. (2023). The encoding and decoding part of the architecture are implemented with convolutional operations, similar to how we implemented the autoencoder. In the bottleneck, the architecture uses a vision transformer (Dosovitskiy et al., 2021).
@@ Appendix C Experiments
Given these datasets, we optimise all models with Adam (Kingma & Ba, 2017) with a batch size of 16. The learning rate is linearly increased from $1\times 10^{-6}$ to $2\times 10^{-4}$ within $5\times 10^{3}$ iterations. Afterwards a cosine scheduler is used to decrease the learning rate again up to $1\times 10^{-6}$ after $10^{5}$ iterations, where the training is stopped. All models are trained without weight decay and other regularisation techniques like dropout, and we select the models that achieve the lowest validation loss. We train the models with mixed precision in bfloat16 and evaluate in single precision. All models are trained on either a NVIDIA RTX5000 GPU (24 GB) or a NVIDIA RTX6000 GPU (48 GB); to train on the RTX5000, we use a batch size of 8 and accumulate the gradient of two iterations. All models are implemented in PyTorch (Paszke et al., 2019) with PyTorch lightning (Falcon et al., 2020) and Hydra (Yadan, 2019). The code is available under https://github.com/cerea-daml/ldm_nextsim.
@@
Both, the reconstructions and the generated data are compared to the testing dataset. For the reconstructions, we perform a one-to-one comparison, while we evaluate the statistics of our generated data. In total, we hence generate $4383$ samples with the diffusion models. Although a smaller number than the $50000$ samples normally used in validating image diffusion models, it results into a fair comparison to the testing dataset. In Appendix E.2, we demonstrate that the dataset size is big enough to see differences in the models.
@@ E.2 Influence of the generated dataset size on the diffusion scores
E.3 Clipping of the generative diffusion model without censored distributions
To avoid unphysical samples during generation, the values from the denoiser are commonly clipped into either fixed (e.g., Ho et al., 2020) or dynamical bounds (Saharia et al., 2022b). However, as shown in Appendix A.1, generative diffusion is built around the assumption of unbounded Gaussian diffusion, and the clipping would violate the Gaussian assumption, which has been applied during training. This could introduce biases into the generation.
In Table 5, we show the results of such a clipped diffusion model in addition to the results of the diffusion models as presented in the main manuscript. For this clipped experiment, we use the pre-trained data-space diffusion model and clip the output of its denoiser into the physical bounds ($\text{SIT}\in[0,\infty)$, $\text{SIC}\in[0,1]$, and $\text{SID}\in[0,1]$). Additionally, we show what happens if we clip the output of the latent diffusion model (LDM) trained without the censored distribution.
Table 5: Comparison of the diffusion models tested in Sec. 3 with diffusion model where the output of the denoiser is clipped into physical bounds or a latent diffusion model where the output of the autoencoder is clipped without a censored distribution.
Model $\mathrm{FAED}$downarrow $\text{RMSE}_{\text{SIE}}$downarrow
Validation 2.38 7.02
Diffusion 1.73 9.66
Diffusion (denoiser clipped) 1.76 12.82
LDM 1.65 11.56
LDM (censored) 1.79 7.32
LDM (output clipped) 1.81 11.04
Including a clipping operator into the denoiser that defines the diffusion model in data space has a negative impact on the physical representation of the sea ice, the RMSE in the sea-ice extent is increased compared to the unclipped diffusion model. While this technique is efficient in image generation, it introduces a bias into the denoiser, which seemingly hurts its performance in our sea-ice generator. Because of biasing the diffusion model, we suspect that clipping introduces a degradation in the cross-correlation between variables, which can have a higher impact on geophysical applications than on image generation.
Clipping the output of a LDM trained with a Gaussian distribution has a neutral impact on the physical representation. Since the model is trained without accounting for clipping during training, the model still produces spurious values for target values where the bounds are reached, only values exceeding the bounds are clipped. For the targets that lay on the bounds, the censored distribution, introduced in Appendix A.2, pushes the output of the LDM towards an extreme exceedance of the bounds, increasing the probability that the values is clipped into the bounds. Hence, censored distributions can unlock an effective incorporation of physical bounds into the training of deep learning algorithms.
This results suggests that we need special methods to properly treat physical bounds in diffusion models that work in data space. We can define a log-barrier that pushes the diffusion process away from the bounds (Fishman et al., 2023). Additionally, we can define a reflected diffusion process, where the Brownian motion is reflected at the bounds (Lou & Ermon, 2023). However, with both methods, we would never generate values that exactly on the bounds. If the result from the diffusion model is further used in downstream applications, e.g., by cycling in a surrogate model (Price et al., 2024; Finn et al., 2024), such small errors can amplify and make the pipeline unstable. Contrastingly, censoring of the autoencoder provides a clean and systematic way to incorporate bounds so that the generated data can lie exactly on the bounds.
E.4 Censoring
@@
E.5 Uncurated data samples
@@
Appendix F Changes to the Version 1
  • We added the comparison in speed to Table 2 indicating that latent diffusion models are indeed much faster than diffusion models in data space.
  • We strengthened the proof-of-concept character of the study, indicating that we see this work as a first step towards the training of generative diffusion for e.g., Arctic-wide surrogate models.
  • We added Appendix E.3 with additional results if we add clipping to the denoiser in data space. While in image generation, this clipping is very important, we find a negative impact on the physical representation in the diffusion model. We additionally provide insides why a censored distribution helps to improve the physical representation.
  • We added the GitHub link to the code.